OpenStudy (anonymous):

how do i prove d/dx cos(x^2)=-2sin(x^2) is true/false?

OpenStudy (anonymous):

Ahhhh headache

OpenStudy (anonymous):

take the derivative using the chain rule and see that you get something different

OpenStudy (anonymous):

the derivative of \(x^2\) is \(2x\) so the correct answer is \[-2x\sin(x^2)\]

OpenStudy (anonymous):

but im confused as hot to start the chain rule...

OpenStudy (anonymous):

how*

OpenStudy (anonymous):

the derivative of \(\cos(x)\) is \(-\sin(x)\) by the chain rule the derivative of \(\cos(f(x))\) is \(-\sin(f(x))\times f'(x)\)

OpenStudy (anonymous):

so you get -sin(x^2) x 2x and we get -2xsin(x^2)?

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

yes yw